# State-insensitive wavelengths for light shifts and photon scattering from Zeeman states

### v1

@@ E.1 Finding $\Delta_{\perp}$
$F=I-\frac{1}{2}:~{}~{}~{}\Delta_{\perp}$ $=\frac{\zeta_{+1}(2F+1)\begin{Bmatrix}J&J^{\prime}&1\\ F&F&I\end{Bmatrix}^{2}}{F(2F+3)\begin{Bmatrix}J&J^{\prime}&1\\ F+1&F&I\end{Bmatrix}^{2}-(2F+1)\begin{Bmatrix}J&J^{\prime}&1\\ F&F&I\end{Bmatrix}^{2}},$ (71g)
$F=I+\frac{1}{2}:~{}~{}~{}\Delta_{\perp}$ $=\frac{\zeta_{-1}(2F+1)\begin{Bmatrix}J&J^{\prime}&1\\ F&F&I\end{Bmatrix}^{2}}{(1-2F)(F+1)\begin{Bmatrix}J&J^{\prime}&1\\ F-1&F&I\end{Bmatrix}^{2}-(2F+1)\begin{Bmatrix}J&J^{\prime}&1\\ F&F&I\end{Bmatrix}^{2}}.$ (71n)
@@ E.2 Finding $\Delta_{\parallel}$
$F=I-\frac{1}{2}:~{}~{}~{}\Delta_{\parallel}$ $=\frac{\zeta_{+1}(4F+2)\begin{Bmatrix}J&J^{\prime}&1\\ F&F&I\end{Bmatrix}^{2}}{2F\begin{Bmatrix}J&J^{\prime}&1\\ F+1&F&I\end{Bmatrix}^{2}-(2F+1)\begin{Bmatrix}J&J^{\prime}&1\\ F&F&I\end{Bmatrix}^{2}},$ (72g)
$F=I+\frac{1}{2}:~{}~{}~{}\Delta_{\parallel}$ $=\frac{\zeta_{-1}(4F+2)\begin{Bmatrix}J&J^{\prime}&1\\ F&F&I\end{Bmatrix}^{2}}{(2F+2)\begin{Bmatrix}J&J^{\prime}&1\\ F-1&F&I\end{Bmatrix}^{2}-(2F+1)\begin{Bmatrix}J&J^{\prime}&1\\ F&F&I\end{Bmatrix}^{2}}$ (72n)
@@ Appendix F Effective Hamiltonian treatment for cavity QED
As in the main text, we consider an atom initialized in a hyperfine state $F$, with $12F+$ 1 magnetic sublevels $\set{-m,-m+1,\dots,m}$ and driven on the $D_{2}$ line by an arbitrary superposition of $\sigma_{\pm}$-polarized light. The atom can scatter into a cavity (or equivalently a waveguide) [see Figure 5]. The cavity has two modes with $\sigma_{\pm}$ polarization and degenerate frequency.
@@
where $\delta=\omega-\omega_{c}$ is the detuning between drive and cavity mode, $\Delta_{F^{\prime}}=\omega-\omega_{F^{\prime}}$ is the detuning of the drive from the excited manifold $F^{\prime}$. We assume that all Zeeman shifts are negligible such that all magnetic sublevels can be considered at the same energy. The interaction Hamiltonian is
$\hat{H}_{\mathrm{int.}}=\sum\limits_{F,F^{\prime}}\left(\frac{\cos\theta~{}\Omega}{2}+g\hat{a}\right)\sum\limits_{m}r^{m}_{F,F^{\prime}}\ket{F^{\prime},m+1}\bra{F,m}+\sum\limits_{F,F^{\prime}}\left(\frac{\sin\theta~{}\Omega}{2}+g\hat{b}\right)\sum\limits_{m}s^{m}_{F,F^{\prime}}\ket{F^{\prime},m-1}\bra{F,m}+\mathrm{H.c.},$ (75)
@@
In the limit of $\Omega_{\pm}/\Delta_{F^{\prime}}\ll 1~{}\forall~{}F^{\prime}$, the atomic saturation is essentially zero, and one can adiabatically eliminate the excited states. This yields an effective Hamiltonian
@@ F.1 Simplifications at the magic detuning
Using these expressions and assuming we operate at a magic detuning, the above Hamiltonian reads

### v2

@@ E.1 Finding $\Delta_{\perp}$
$F=I-\frac{1}{2}:~{}~{}~{}\Delta_{\perp}$ $=\frac{\zeta_{+1}(2F+1)\begin{Bmatrix}J&J^{\prime}&1\\ F&F&I\end{Bmatrix}^{2}}{F(2F+3)\begin{Bmatrix}J&J^{\prime}&1\\ F+1&F&I\end{Bmatrix}^{2}-(2F+1)\begin{Bmatrix}J&J^{\prime}&1\\ F&F&I\end{Bmatrix}^{2}},$ (69a)
$F=I+\frac{1}{2}:~{}~{}~{}\Delta_{\perp}$ $=\frac{\zeta_{-1}(2F+1)\begin{Bmatrix}J&J^{\prime}&1\\ F&F&I\end{Bmatrix}^{2}}{(1-2F)(F+1)\begin{Bmatrix}J&J^{\prime}&1\\ F-1&F&I\end{Bmatrix}^{2}-(2F+1)\begin{Bmatrix}J&J^{\prime}&1\\ F&F&I\end{Bmatrix}^{2}}.$ (69b)
@@ E.2 Finding $\Delta_{\parallel}$
$F=I-\frac{1}{2}:~{}~{}~{}\Delta_{\parallel}$ $=\frac{\zeta_{+1}(4F+2)\begin{Bmatrix}J&J^{\prime}&1\\ F&F&I\end{Bmatrix}^{2}}{2F\begin{Bmatrix}J&J^{\prime}&1\\ F+1&F&I\end{Bmatrix}^{2}-(2F+1)\begin{Bmatrix}J&J^{\prime}&1\\ F&F&I\end{Bmatrix}^{2}},$ (70a)
$F=I+\frac{1}{2}:~{}~{}~{}\Delta_{\parallel}$ $=\frac{\zeta_{-1}(4F+2)\begin{Bmatrix}J&J^{\prime}&1\\ F&F&I\end{Bmatrix}^{2}}{(2F+2)\begin{Bmatrix}J&J^{\prime}&1\\ F-1&F&I\end{Bmatrix}^{2}-(2F+1)\begin{Bmatrix}J&J^{\prime}&1\\ F&F&I\end{Bmatrix}^{2}}$ (70b)
@@ Appendix F Effective Hamiltonian treatment for cavity QED
As in the main text, we consider an atom initialized in a hyperfine state $F$, with $12F+$ 1 Zeeman states $\set{-m,-m+1,\dots,m}$ and driven on the $D_{2}$ line by an arbitrary superposition of $\sigma_{\pm}$-polarized light. The atom can scatter into a cavity (or equivalently a waveguide) [see Figure 5]. The cavity has two modes with $\sigma_{\pm}$ polarization and degenerate frequency.
@@
where $\delta=\omega-\omega_{c}$ is the detuning between drive and cavity mode, $\Delta_{F^{\prime}}=\omega-\omega_{F^{\prime}}$ is the detuning of the drive from the excited manifold $F^{\prime}$. We assume that all Zeeman shifts are negligible such that all Zeeman states can be considered at the same energy. The interaction Hamiltonian is
$\hat{H}_{\mathrm{int.}}=\sum\limits_{F,F^{\prime}}\left(\frac{\cos\theta~{}\Omega}{2}+g\hat{a}\right)\sum\limits_{m}r_{m,F,F^{\prime}}\ket{F^{\prime},m+1}\bra{F,m}+\sum\limits_{F,F^{\prime}}\left(\frac{\sin\theta~{}\Omega}{2}+g\hat{b}\right)\sum\limits_{m}s_{m,F,F^{\prime}}\ket{F^{\prime},m-1}\bra{F,m}+\mathrm{H.c.},$ (73)
@@
In the limit of $\Omega/\Delta_{F^{\prime}}\ll 1~{}\forall~{}F^{\prime}$, the atomic saturation is essentially zero, and one can adiabatically eliminate the excited states. This yields an effective Hamiltonian
@@ F.1 Simplifications at the magic detuning
Using these expressions and assuming we operate at a magic detuning, the above Hamiltonian reads as